OpenStudy (anonymous):

The distance d when a spring is stretched by a hanging object varies directly as the weight w of the object. If the distance is 61 cm when the weight is 4 kg, what is the distance when the weight is 9 kg?

OpenStudy (anonymous):

9/4=x/61 9*61/4=x can you solve after this?

OpenStudy (anonymous):

let me try

OpenStudy (anonymous):

X=549/4?

OpenStudy (anonymous):

no let me take that back thats not it

OpenStudy (anonymous):

yeah that's it. i think you need to get an answer in decimals or something.

OpenStudy (anonymous):

oh really, ok thanks ! for the help